Sound boxes for tuning forks are essentially pipes with one end open and one end closed. The energy of the vibrating fork builds up a standing pressure wave in the sound box. Some of the sound leaks out and makes the tuning fork more audible than it would be without the sound box. Design a box for a tuning fork at 440 Hz.
Answer #1

The fundamental frequency of a closed air column =

f=\frac{v_{sound}}{4L}=440\\ \\\Rightarrow L=\frac{340}{4*440}=0.18m

So an 18cm closed end tube will be perfect for a sound box.

Note: in practice, 0.6 times the radius of the pipe should be added to the open end to get the proper length.